# MF-GLaM: A multifidelity stochastic emulator using generalized lambda models

### v1

@@ 1 Introduction
Emulating stochastic simulators can be computationally demanding due to the large number of model evaluations required to estimate full response distributions accurately. When emulating expensive high-fidelity (HF) models, the associated computational cost can be prohibitive. In scenarios where stochastic simulators of different fidelities exist, we can take advantage of the lower-fidelity (LF) model to emulate the higher fidelity one. This is the core idea behind multifidelity surrogate modeling. Multifidelity surrogate models (MFSMs) combine data from multiple sources of different fidelity into a single surrogate model, typically enhancing a limited, computationally expensive HF dataset with more extensive and less expensive LF ones. Over the past two decades, multifidelity surrogate modeling has been extensively studied in the context of deterministic simulators, leading to a variety of successful approaches, including Gaussian process (GP)-based methods (Kennedy and O’Hagan, 2000; Le Gratiet and Garnier, 2014), deep GP-based approaches (Cutajar et al., 2018; Hebbal et al., 2021), PCE-based approaches (Ng and Eldred, 2012; Palar et al., 2016), and neural network-based frameworks (Zhang et al., 2022; Conti et al., 2023). However, not all multifidelity approaches build a single combined surrogate model. In probabilistic UQ, for example, fidelity hierarchies can be directly exploited via Monte Carlo methods, such as multilevel Monte Carlo (Giles, 2008), approximate control-variate estimators (Gorodetsky et al., 2020), multifidelity importance sampling (Dubourg and Sudret, 2014; Peherstorfer et al., 2016), and control variates importance sampling (Chakroborty et al., 2024). Likewise, in interval-based UQ, multilevel interval-analysis methods have been developed (Callens et al., 2022).
@@
However, constructing multifidelity surrogate models to emulate inherently stochastic simulators, where variability arises intrinsically and the homoscedasticity assumption is relaxed, has received comparatively little attention until recently. Perdikaris et al. (2015) propose a recursive co-Kriging framework to integrate multifidelity information both in the models and in the probability space for design under uncertainty, focusing on estimating the mean or other statistical quantities of interest of a HF simulator. The first attempt to apply multifidelity modeling to correlated stochastic processes for learning the conditional distribution of the HF process is proposed by Yang and Perdikaris (2019), who propose a surrogate model by performing variational inference using adversarial learning. Moreover, Stroh et al. (2021) introduce an adaptive sampling scheme based on GP modeling, applicable to both deterministic and stochastic multifidelity models. In addition, Bae et al. (2020) develop a nondeterministic localized-Galerkin approach extending the nondeterministic Kriging (Bae et al., 2019) that can handle both aleatoric and epistemic uncertainties. Tao et al. (2025) propose an MF co-Kriging approach for stochastic simulators with heteroscedastic noise, while Konomi et al. (2023) propose a Bayesian latent-variable co-Kriging model for handling quality-flagged observations in remote sensing applications. In the relevant context of uncertainty quantification, Reuter et al. (2024) extend the approximate control variate framework (Gorodetsky et al., 2020) for multifidelity UQ in stochastic simulators, yet without constructing surrogate models that emulate the entire output distribution. Although these methods show significant progress, certain challenges remain. Several approaches focus on specific summary statistics, such as the mean, variance, or selected quantiles, rather than modeling the full conditional response distribution of high-fidelity stochastic emulators. In addition, a generally applicable and computationally efficient multifidelity stochastic emulator that can flexibly capture highly non-Gaussian response distributions from limited stochastic data has yet to be established.
@@
The remainder of this paper is organized as follows. Firstly, we recall the GLaM methodology in the single-fidelity setting in Section 2, along with the relevant theory. Next, we introduce the MF-GLaM framework and evaluate its performance on three numerical examples, in Sections 3 and 4, respectively. Finally, in Section 5, we draw conclusions and discuss prospects for future research.
@@ 2 Generalized lambda models (GLaMs) in single fidelity
When applied to stochastic emulation, each of the four parameters $\boldsymbol{\lambda}$ in Eq. 2 is modeled as a function of $\boldsymbol{x}$, allowing the shape of the distribution to vary throughout the input space, a property necessary to address heteroskedasticity in the model response. To achieve this, Zhu and Sudret (2021) construct a functional approximation of each component of $\boldsymbol{\lambda}(\boldsymbol{x})$ using polynomial chaos expansions. The latter is a surrogate modeling technique that approximates the response of a (deterministic) model with finite variance through its spectral representation on an orthogonal polynomial basis (Xiu and Karniadakis, 2002; Ghanem and Spanos, 1991). In practice, this series is truncated to a finite number of terms. Consider a random vector $\boldsymbol{X}\in\mathbb{R}^{M}$ with independent components and joint PDF $f_{\boldsymbol{X}}(\boldsymbol{x})=\prod_{i=1}^{M}f_{{X}_{i}}(x_{i}),$ where $f_{{X}_{i}}$ is the marginal PDF of the random variable ${X}_{i}$. The truncated PCE of a model ${\mathcal{M}}(\boldsymbol{x})$ is expressed as
~$\widetilde{\mathcal{M}}^{\text{PC}}\left(\boldsymbol{x}\right)=\sum_{\boldsymbol{\alpha}\in{\mathcal{A}}}c_{\boldsymbol{\alpha}}{\Psi}_{\boldsymbol{\alpha}}\left(\boldsymbol{x}\right),$ (9)
@@
where $f^{\text{GLD}}$ is the PDF of the GLD, as defined in Eq. 4.
The truncation sets ${\mathcal{A}}_{1}$ and ${\mathcal{A}}_{2}$ corresponding to $\lambda_{1}^{\text{PC}}$ and $\lambda_{2}^{\text{PC}}$, respectively, are obtained using a modified feasible generalized least-squares approach (Zhu and Sudret, 2021). More precisely, the mean and variance functions of the model response are each expressed as PCEs and are fitted alternatively and iteratively using the sparse solver hybrid least-angle regression (Blatman and Sudret, 2011). These fitted PCEs provide the truncation sets for $\lambda_{1}^{\text{PC}}$ and $\lambda_{2}^{\text{PC}}$, as well as starting points $\boldsymbol{c}_{1,0},\boldsymbol{c}_{2,0}$ for the optimization in Eq. 13.
@@ 3.1 Introduction
We also use PCE to model the discrepancy functions $\boldsymbol{\delta}$ in Eq. 19. Hence, $\boldsymbol{\lambda}^{\text{MF}}$ is written as:
@@ 3.2 Estimation of the model parameters
From Eqs. 20, 21, 22 and 23, we observe that some of the coefficients in $\boldsymbol{\theta}$ appear in both the LF GLaM and the MF-GLaM, namely, the coefficients $\boldsymbol{c}$. On the other hand, the coefficients $\boldsymbol{d}$ appear only in the MF-GLaM. Our goal is thus to jointly estimate all unknown coefficients $\boldsymbol{\theta}$ using both the HF and LF data simultaneously. To achieve this, we employ a maximum likelihood estimation approach. More precisely, we derive such an estimator by minimizing the Kullback-Leibler divergence between the true underlying joint distribution of the HF and LF inputs and outputs, and the distribution approximated by GLaMs. To do this, we will need to delve deeper into the nature of multifidelity modeling for stochastic models.
@@
[…]
where $\mathbb{P}(S=L)=p$, $\mathbb{P}(S=H)=1-p$, and $f_{\boldsymbol{X}}$ is the PDF of $\boldsymbol{X}$. Here, $p$ represents the prior probability that the label $S$ takes the value $L$, corresponding to selecting the LF model, and captures the relative weight of sampling LF data compared to the HF data. Moreover, $f_{Y_{L}\mid\boldsymbol{X}}\bigl{(}y_{L}\mid\boldsymbol{x}\bigr{)}$ and $f_{Y_{H}\mid\boldsymbol{X}}\bigl{(}y_{H}\mid\boldsymbol{x}\bigr{)}$ denote the conditional densities of $Y_{L}$ and $Y_{H}$, respectively, given their input variables. We note that each conditional density depends only on the subset of inputs that the specific model uses. Therefore, in Eq. 27 we have ignored the irrelevant input variables when expressing the conditional distributions, e.g., $f_{Y_{L}\mid\boldsymbol{X}}(y_{L}\mid\boldsymbol{x})$.
A GLaM surrogate for the LF response provides an approximation $f_{Y_{L}\mid\boldsymbol{X},\mathrm{GLaM}}\bigl{(}y_{L}\mid\boldsymbol{\lambda}^{\text{L, PC}}(\boldsymbol{x};\boldsymbol{c})\bigr{)}$ to the conditional PDF $f_{Y_{L}\mid\boldsymbol{X}}\bigl{(}y_{L}\mid\boldsymbol{x}\bigr{)}$. For brevity, the GLaM approximation to the LF conditional PDF is denoted as $f_{\mathrm{GLaM}}^{\text{L}}\bigl{(}y_{L}\mid\boldsymbol{x};\boldsymbol{c}\bigr{)}$. Moreover, the multifidelity GLaM surrogate provides an approximation $f_{Y_{H}\mid\boldsymbol{X},\mathrm{GLaM}}\bigl{(}y_{H}\mid\boldsymbol{\lambda}^{\text{H, PC}}(\boldsymbol{x};\boldsymbol{c},\boldsymbol{d})\bigr{)}$, denoted as $f_{\mathrm{GLaM}}^{\text{MF}}\bigl{(}y_{H}\mid\boldsymbol{x};\boldsymbol{c},\boldsymbol{d}\bigr{)}$, to the HF conditional PDF $f_{Y_{H}\mid\boldsymbol{X}}\bigl{(}y_{H}\mid\boldsymbol{x}\bigr{)}$. Thus, the GLaM approximation to the joint PDF in Eq. 27 is obtained as follows:
@@
The detailed derivation to obtain this estimator is provided in Appendix A.
In common multifidelity settings, the HF data are typically available in much smaller quantities. Relying solely on a limited HF dataset (i.e., $p=0$) to train a model could lead to overfitting, since the model does not have enough information to generalize reliably. Thus, the LF data can act as a form of “regularization”. A practical compromise, in the absence of domain-specific guidance, is to assign a neutral weight $p=0.5$ between LF and HF in the likelihood function, ensuring that neither fidelity dominates simply because of its dataset size. Under this assumption, from Eq. 29, the log-likelihood to be maximized with respect to the parameters $\boldsymbol{\theta}$ is given by:
@@ 3.3 Basis selection and fitting procedure
In this section, we present a method for determining suitable truncation schemes $\boldsymbol{{\mathcal{A}}}^{\text{L}},\boldsymbol{{\mathcal{A}}}^{\delta}$ for the expansions of $\{\lambda_{i}^{\text{MF}}:i=1,...,4\}$ in Eqs. 22 and 23, followed by the practical fitting procedure used to obtain the corresponding coefficients according to the optimization in Eq. 31. This also includes selecting good initial values for $\boldsymbol{\theta}$ in order to facilitate convergence of the optimization.
The overall idea is to first train a GLaM using only the LF data to define both the starting points for the MF-GLaM coefficients and the basis functions that correspond to the LF expansion. Once these have been identified, the full HF and LF datasets are used to jointly maximize Eq. 30, thereby estimating the final coefficients $\boldsymbol{c}$ and $\boldsymbol{d}$. We apply this procedure for multiple different configurations of the discrepancy truncation sets, with the best one identified via a modified Bayesian information criterion, described subsequently.
More precisely, starting from the LF GLaM, we obtain both the truncation sets $\{{\mathcal{A}}_{i}^{\text{L}}:i=1,...,4\}$ and initial coefficients $\{\boldsymbol{c}_{i,0}:i=1,...,4\}$. We now address the truncation sets $\{{\mathcal{A}}_{i}^{\delta}:i=1,...,4\}$ and initial coefficients $\{\boldsymbol{d}_{i,0}:i=1,...,4\}$ associated with the discrepancy PCEs, starting by considering a simplification regarding the relation between HF and LF GLaM shapes. As discussed in Section 2, the shape of the PDF of a GLaM is primarily determined by the two parameters $\lambda_{3}$ and $\lambda_{4}$. In practice, we often find that the shape of the LF response PDF closely matches the corresponding HF shape. Therefore, we set the discrepancies $\delta_{3}^{\text{PC}},\delta_{4}^{\text{PC}}$ to zero ($\delta_{3}^{\mathrm{PC}}=\delta_{4}^{\mathrm{PC}}=0$) in Eq. 22, which gives us:
@@
The expressions for $\lambda_{1}$ and $\lambda_{2}$ remain as in Eq. 22 and Eq. 23, respectively. To select the truncation sets $\{{\mathcal{A}}_{i}^{\delta}:i=1,2\}$, we examine several candidate degrees $\boldsymbol{p}_{1,\delta}$ and $\boldsymbol{p}_{2,\delta}$, and q-norms $\boldsymbol{q}_{1,\delta}$ and $\boldsymbol{q}_{2,\delta}$. For each unique combination $\{(p_{1,\delta}^{(j)},q_{1,\delta}^{(k)},p_{2,\delta}^{(l)},q_{2,\delta}^{(m)})\}$, we initialize $\boldsymbol{d}_{i,0}\!=0$ (for $i=1,2$) and fit an MF-GLaM on all HF and LF data. Among the resulting models, we select the one that minimizes the following BIC score:
[…]
where $\ell(\boldsymbol{\theta};{\mathcal{S}},\boldsymbol{{\mathcal{X}}},{\mathcal{Y}}_{H},{\mathcal{Y}}_{L})$ is the log-likelihood in Eq. 30 and $n_{\boldsymbol{\theta}}$ is the total number of parameters in the set $\boldsymbol{\theta}$. A full Bayesian derivation of the MF-BIC score in Eq. 33 is given in Appendix B.
The optimization strategy we use to obtain the parameters $\boldsymbol{\theta}$ that maximize the log-likelihood in Eq. 30 is the same as the one described by Zhu and Sudret (2020, 2021). More precisely, the derivative-based trust-region method (Steihaug, 1983) is applied first to solve the unconstrained problem. If constraints are violated (due to the dependence of the GLD support on the distribution parameters) we switch to the constrained (1+1)-CMA-ES algorithm (Arnold and Hansen, 2012) to obtain the final solution.

### v2

# 11affiliationtext: Chair of Risk, Safety and Uncertainty Quantification, ETH Zürich, Switzerland22affiliationtext: Laboratory of Signals and Systems, CentraleSupélec, Université Paris-Saclay

@@ 1 Introduction
Emulating stochastic simulators can be computationally demanding due to the large number of model evaluations required to estimate full response distributions accurately. When emulating expensive high-fidelity (HF) models, the associated computational cost can be prohibitive. In scenarios where stochastic simulators of different fidelities exist, we can take advantage of the lower-fidelity (LF) model to emulate the higher fidelity one. This is the core idea behind multifidelity surrogate modeling. Multifidelity surrogate models (MFSMs) combine data from multiple sources of different fidelity into a single surrogate model, typically enhancing a limited, computationally expensive HF dataset with more extensive and less expensive LF ones. Over the past two decades, multifidelity surrogate modeling has been extensively studied in the context of deterministic simulators, leading to a variety of successful approaches, including Gaussian process (GP)-based methods (Kennedy and O’Hagan, 2000; Le Gratiet and Garnier, 2014), deep GP-based approaches (Cutajar et al., 2018; Hebbal et al., 2021), PCE-based approaches (Ng and Eldred, 2012; Palar et al., 2016), and neural network-based frameworks (Zhang et al., 2022; Conti et al., 2023). However, not all multifidelity approaches build a single combined surrogate model. In probabilistic UQ, for example, fidelity hierarchies can be directly exploited via Monte Carlo methods, such as multilevel Monte Carlo (Giles, 2008), approximate control‐variate estimators (Gorodetsky et al., 2020), multifidelity importance sampling (Dubourg and Sudret, 2014; Peherstorfer et al., 2016), and control variates importance sampling (Chakroborty et al., 2024). Likewise, in interval‐based UQ, multilevel interval‐analysis methods have been developed (Callens et al., 2022).
@@
However, constructing multifidelity surrogate models to emulate inherently stochastic simulators, where variability arises intrinsically and the homoscedasticity assumption is relaxed, has received comparatively little attention until recently. Perdikaris et al. (2015) propose a recursive co-Kriging framework to integrate multifidelity information both in the models and in the probability space for design under uncertainty, focusing on estimating the mean or other statistical quantities of interest of a HF simulator. The first attempt to apply multifidelity modeling to correlated stochastic processes for learning the conditional distribution of the HF process is proposed by Yang and Perdikaris (2019), who propose a surrogate model by performing variational inference using adversarial learning. Moreover, Stroh et al. (2021) introduce an adaptive sampling scheme based on GP modeling, applicable to both deterministic and stochastic multifidelity models. In addition, Bae et al. (2020) develop a nondeterministic localized‑Galerkin approach extending the nondeterministic Kriging (Bae et al., 2019) that can handle both aleatoric and epistemic uncertainties. Tao et al. (2025) propose an MF co-Kriging approach for stochastic simulators with heteroscedastic noise, while Konomi et al. (2023) propose a Bayesian latent‑variable co‑Kriging model for handling quality‑flagged observations in remote sensing applications. In the relevant context of uncertainty quantification, Reuter et al. (2024) extend the approximate control variate framework (Gorodetsky et al., 2020) for multifidelity UQ in stochastic simulators, yet without constructing surrogate models that emulate the entire output distribution. Although these methods show significant progress, certain challenges remain. Several approaches focus on specific summary statistics, such as the mean, variance, or selected quantiles, rather than modeling the full conditional response distribution of high-fidelity stochastic emulators. In addition, a generally applicable and computationally efficient multifidelity stochastic emulator that can flexibly capture highly non-Gaussian response distributions from limited stochastic data has yet to be established.
@@
The remainder of this paper is organized as follows. Firstly, we recall the GLaM methodology in the single-fidelity setting in Section˜2, along with the relevant theory. Next, we introduce the MF-GLaM framework and evaluate its performance on three numerical examples, in Sections˜3 and 4, respectively. Finally, in Section˜5, we draw conclusions and discuss prospects for future research.
@@ 2 Generalized lambda models (GLaMs) in single fidelity
When applied to stochastic emulation, each of the four parameters $\boldsymbol{\lambda}$ in Eq.˜2 is modeled as a function of $\boldsymbol{x}$, allowing the shape of the distribution to vary throughout the input space, a property necessary to address heteroskedasticity in the model response. To achieve this, Zhu and Sudret (2021) construct a functional approximation of each component of $\boldsymbol{\lambda}(\boldsymbol{x})$ using polynomial chaos expansions. The latter is a surrogate modeling technique that approximates the response of a (deterministic) model with finite variance through its spectral representation on an orthogonal polynomial basis (Xiu and Karniadakis, 2002; Ghanem and Spanos, 1991). In practice, this series is truncated to a finite number of terms. Consider a random vector $\boldsymbol{X}\in\mathbb{R}^{M}$ with independent components and joint PDF $f_{\boldsymbol{X}}(\boldsymbol{x})=\prod_{i=1}^{M}f_{{X}_{i}}(x_{i}),$ where $f_{{X}_{i}}$ is the marginal PDF of the random variable ${X}_{i}$. The truncated PCE of a model ${\mathcal{M}}(\boldsymbol{x})$ is expressed as
$\widetilde{\mathcal{M}}^{\text{PC}}\left(\boldsymbol{x}\right)=\sum_{\boldsymbol{\alpha}\in{\mathcal{A}}}c_{\boldsymbol{\alpha}}{\Psi}_{\boldsymbol{\alpha}}\left(\boldsymbol{x}\right),$ (9)
@@
where $f^{\text{GLD}}$ is the PDF of the GLD, as defined in Eq.˜4.
The truncation sets ${\mathcal{A}}_{1}$ and ${\mathcal{A}}_{2}$ corresponding to $\lambda_{1}^{\text{PC}}$ and $\lambda_{2}^{\text{PC}}$, respectively, are obtained using a modified feasible generalized least-squares approach (Zhu and Sudret, 2021). More precisely, the mean and variance functions of the model response are each expressed as PCEs and are fitted alternatively and iteratively using the sparse solver hybrid least-angle regression (Blatman and Sudret, 2011). These fitted PCEs provide the truncation sets for $\lambda_{1}^{\text{PC}}$ and $\lambda_{2}^{\text{PC}}$, as well as starting points $\boldsymbol{c}_{1,0},\boldsymbol{c}_{2,0}$ for the optimization in Eq.˜13.
@@ 3.1 Introduction
We also use PCE to model the discrepancy functions $\boldsymbol{\delta}$ in Eq.˜19. Hence, $\boldsymbol{\lambda}^{\text{MF}}$ is written as:
@@ 3.2 Estimation of the model parameters
From Eqs.˜20, 21, 22 and 23, we observe that some of the coefficients in $\boldsymbol{\theta}$ appear in both the LF GLaM and the MF-GLaM, namely, the coefficients $\boldsymbol{c}$. On the other hand, the coefficients $\boldsymbol{d}$ appear only in the MF-GLaM. Our goal is thus to jointly estimate all unknown coefficients $\boldsymbol{\theta}$ using both the HF and LF data simultaneously. To achieve this, we employ a maximum likelihood estimation approach. More precisely, we derive such an estimator by minimizing the Kullback-Leibler divergence between the true underlying joint distribution of the HF and LF inputs and outputs, and the distribution approximated by GLaMs. To do this, we will need to delve deeper into the nature of multifidelity modeling for stochastic models.
@@
$f_{0}\bigl(s,\boldsymbol{x},y_{H},y_{L}\bigr)\;=\;p\,\mathbbm{1}_{\{s=L\}}\,f_{\boldsymbol{X}}(\boldsymbol{x})\,f_{Y_{L}\mid\boldsymbol{X}}\bigl(y_{L}\mid\boldsymbol{x}\bigr)\;+\;(1-p)\,\mathbbm{1}_{\{s=H\}}\,f_{\boldsymbol{X}}(\boldsymbol{x})\,f_{Y_{H}\mid\boldsymbol{X}}\bigl(y_{H}\mid\boldsymbol{x}\bigr),$ (27)
where $\mathbb{P}(S=L)=p$, $\mathbb{P}(S=H)=1-p$, and $f_{\boldsymbol{X}}$ is the PDF of $\boldsymbol{X}$. Here, $p$ represents the prior probability that the label $S$ takes the value $L$, corresponding to selecting the LF model, and captures the relative weight of sampling LF data compared to the HF data. Moreover, $f_{Y_{L}\mid\boldsymbol{X}}\bigl(y_{L}\mid\boldsymbol{x}\bigr)$ and $f_{Y_{H}\mid\boldsymbol{X}}\bigl(y_{H}\mid\boldsymbol{x}\bigr)$ denote the conditional densities of $Y_{L}$ and $Y_{H}$, respectively, given their input variables. We note that each conditional density depends only on the subset of inputs that the specific model uses. Therefore, in Eq.˜27 we have ignored the irrelevant input variables when expressing the conditional distributions, e.g., $f_{Y_{L}\mid\boldsymbol{X}}(y_{L}\mid\boldsymbol{x})$.
A GLaM surrogate for the LF response provides an approximation $f_{Y_{L}\mid\boldsymbol{X},\mathrm{GLaM}}\bigl(y_{L}\mid\boldsymbol{\lambda}^{\text{L, PC}}(\boldsymbol{x};\boldsymbol{c})\bigr)$ to the conditional PDF $f_{Y_{L}\mid\boldsymbol{X}}\bigl(y_{L}\mid\boldsymbol{x}\bigr)$. For brevity, the GLaM approximation to the LF conditional PDF is denoted as $f_{\mathrm{GLaM}}^{\text{L}}\bigl(y_{L}\mid\boldsymbol{x};\boldsymbol{c}\bigr)$. Moreover, the multifidelity GLaM surrogate provides an approximation $f_{Y_{H}\mid\boldsymbol{X},\mathrm{GLaM}}\bigl(y_{H}\mid\boldsymbol{\lambda}^{\text{H, PC}}(\boldsymbol{x};\boldsymbol{c},\boldsymbol{d})\bigr)$, denoted as $f_{\mathrm{GLaM}}^{\text{MF}}\bigl(y_{H}\mid\boldsymbol{x};\boldsymbol{c},\boldsymbol{d}\bigr)$, to the HF conditional PDF $f_{Y_{H}\mid\boldsymbol{X}}\bigl(y_{H}\mid\boldsymbol{x}\bigr)$. Thus, the GLaM approximation to the joint PDF in Eq.˜27 is obtained as follows:
@@
The detailed derivation to obtain this estimator is provided in Appendix˜A.
In common multifidelity settings, the HF data are typically available in much smaller quantities. Relying solely on a limited HF dataset (i.e., $p=0$) to train a model could lead to overfitting, since the model does not have enough information to generalize reliably. Thus, the LF data can act as a form of “regularization”. A practical compromise, in the absence of domain-specific guidance, is to assign a neutral weight $p=0.5$ between LF and HF in the likelihood function, ensuring that neither fidelity dominates simply because of its dataset size. Under this assumption, from Eq.˜29, the log-likelihood to be maximized with respect to the parameters $\boldsymbol{\theta}$ is given by:
@@ 3.3 Basis selection and fitting procedure
In this section, we present a method for determining suitable truncation schemes $\boldsymbol{{\mathcal{A}}}^{\text{L}},\boldsymbol{{\mathcal{A}}}^{\delta}$ for the expansions of $\{\lambda_{i}^{\text{MF}}:i=1,...,4\}$ in Eqs.˜22 and 23, followed by the practical fitting procedure used to obtain the corresponding coefficients according to the optimization in Eq.˜31. This also includes selecting good initial values for $\boldsymbol{\theta}$ in order to facilitate convergence of the optimization.
The overall idea is to first train a GLaM using only the LF data to define both the starting points for the MF-GLaM coefficients and the basis functions that correspond to the LF expansion. Once these have been identified, the full HF and LF datasets are used to jointly maximize Eq.˜30, thereby estimating the final coefficients $\boldsymbol{c}$ and $\boldsymbol{d}$. We apply this procedure for multiple different configurations of the discrepancy truncation sets, with the best one identified via a modified Bayesian information criterion, described subsequently.
More precisely, starting from the LF GLaM, we obtain both the truncation sets $\{{\mathcal{A}}_{i}^{\text{L}}:i=1,...,4\}$ and initial coefficients $\{\boldsymbol{c}_{i,0}:i=1,...,4\}$. We now address the truncation sets $\{{\mathcal{A}}_{i}^{\delta}:i=1,...,4\}$ and initial coefficients $\{\boldsymbol{d}_{i,0}:i=1,...,4\}$ associated with the discrepancy PCEs, starting by considering a simplification regarding the relation between HF and LF GLaM shapes. As discussed in Section˜2, the shape of the PDF of a GLaM is primarily determined by the two parameters $\lambda_{3}$ and $\lambda_{4}$. In practice, we often find that the shape of the LF response PDF closely matches the corresponding HF shape. Therefore, we set the discrepancies $\delta_{3}^{\text{PC}},\delta_{4}^{\text{PC}}$ to zero ($\delta_{3}^{\mathrm{PC}}=\delta_{4}^{\mathrm{PC}}=0$) in Eq.˜22, which gives us:
@@
The expressions for $\lambda_{1}$ and $\lambda_{2}$ remain as in Eq.˜22 and Eq.˜23, respectively. To select the truncation sets $\{{\mathcal{A}}_{i}^{\delta}:i=1,2\}$, we examine several candidate degrees $\boldsymbol{p}_{1,\delta}$ and $\boldsymbol{p}_{2,\delta}$, and q-norms $\boldsymbol{q}_{1,\delta}$ and $\boldsymbol{q}_{2,\delta}$. For each unique combination $\{(p_{1,\delta}^{(j)},q_{1,\delta}^{(k)},p_{2,\delta}^{(l)},q_{2,\delta}^{(m)})\}$, we initialize $\boldsymbol{d}_{i,0}\!=0$ (for $i=1,2$) and fit an MF-GLaM on all HF and LF data. Among the resulting models, we select the one that minimizes the following BIC score:
$\text{MF-BIC}=-2\,\ell(\boldsymbol{\theta};{\mathcal{S}},\boldsymbol{{\mathcal{X}}},{\mathcal{Y}}_{H},{\mathcal{Y}}_{L})\,+\,n_{\boldsymbol{\theta}}\log\left(\frac{N_{\text{L}}+N_{\text{H}}}{2}\right).$ (33)
where $\ell(\boldsymbol{\theta};{\mathcal{S}},\boldsymbol{{\mathcal{X}}},{\mathcal{Y}}_{H},{\mathcal{Y}}_{L})$ is the log-likelihood in Eq.˜30 and $n_{\boldsymbol{\theta}}$ is the total number of parameters in the set $\boldsymbol{\theta}$. A full Bayesian derivation of the MF-BIC score in Eq.˜33 is given in Appendix˜B.
The optimization strategy we use to obtain the parameters $\boldsymbol{\theta}$ that maximize the log-likelihood in Eq.˜30 is the same as the one described by Zhu and Sudret (2020, 2021). More precisely, the derivative-based trust-region method (Steihaug, 1983) is applied first to solve the unconstrained problem. If constraints are violated (due to the dependence of the GLD support on the distribution parameters) we switch to the constrained (1+1)-CMA-ES algorithm (Arnold and Hansen, 2012) to obtain the final solution.